I am HORRIBLE with percents, so if you can please help, please and thanks!
Thirty-three pints of acid solution contain 3 pints of pure acid. Find the number of pints of pure acid that must be added to the 33 pints so that the resulting solution is 25% pure acid.

Answer by stanbon(75887)

Thirty-three pints of acid solution contain 3 pints of pure acid. Find the number of pints of pure acid that must be added to the 33 pints so that the resulting solution is 25% pure acid.
-----
Equation:
acid + acid = acid
3 + x = 0.25(33+x)
----
Multiply thru by 100 to get:
300 + 100x = 25*33 + 25x
75x = 525
x = 11 pints (amt. of pure acid needed)
===============
Cheers,
Stan H.
==========